Find dy/dx and d²y/dx² without eliminating the parameter.


a.) x= acosh(t) , y= bsinh(t)

b.) x= acos(t) , y= bsin(t)


1
Expert's answer
2021-01-17T17:15:09-0500

To find "\frac{dy}{dx}" , "\frac{d^2y}{dx^2}" will be used formula "\frac{dy}{dx}=\frac{dy}{dt}/\frac{dx}{dt}" . Such formula are takeen the standard chain rule 

"\frac{dy}{dt} =\frac{ dy}{dx }\frac{ dx}{dt}"

a) Let's take derivatives of the "x" and "y" with respect to the independent variable "t"

"x=a\cosh{t}"

"y=b\sinh{t}"

"\frac{dx}{dt}=a\sinh{t}"

"\frac{dy}{dt}=b\cosh{t}"

Substitute derivatives into formula "\frac{dy}{dx}=\frac{dy}{dt}/\frac{dx}{dt}" :

"\frac{dy}{dx}=\frac{b\cosh{t}}{a\sinh{t}}=\frac{b}{a}\coth{t}"

To find "\frac{d^2y}{dx^2}" , firstly take the derivative of the function for "\frac{dy}{dx}" (as "y'" ) :

"y'=\frac{d}{dt}(\frac{dy}{dx})=\frac{b}{a}(1-\coth^2{t})=-\frac{b}{a}\sinh^{-2}{t}"

"\\frac{d^2y}{dx^2}=\\frac{dy'}{\\frac{dx}{dt}}= \\frac{b}{a}(1-\\coth^2{t})\/(a\\sinh{t})=-\\frac{b}{a^2}\\sinh^3{t}\n\u200b"

b) Let's take derivatives of the "x" and "y" with respect to the independent variable "t"

"x=a\cos{t}"

"y=b\sin{t}"

"\frac{dx}{dt}=-a\sin{t}"

"\frac{dy}{dt}=b\cos{t}"

Substitute derivatives into formula "\frac{dy}{dx}=\frac{dy}{dt}/\frac{dx}{dt}" :

"\frac{dy}{dx}=-\frac{b\cos{t}}{a\sin{t}}=-\frac{b}{a}\cot{t}"

To find "\frac{d^2y}{dx^2}" , firstly take the derivative of the function for "\frac{dy}{dx}" :

"y'=\frac{d}{dt}(\frac{dy}{dx})=-\frac{b}{a}\csc^2{t}"

"\\frac{d^2y}{dx^2}=\\frac{dy'}{\\frac{dx}{dt}}= -\\frac{b}{a}\\csc^2{t}\/(-a\\sin{t})=\\frac{b}{a^2}\\sin^3{t}\n\u200b"